OpenStudy (anonymous):

THE SUM OF TWO CONSECUTIVES INTEGERS IS 59. WRITE AN EQUATION THAT MODELS THIS SITUATION AND FIND THE VALUES OF THE TWO INTEGERS TCARROLL010 PLEASE HELP ME

OpenStudy (anonymous):

59=x+(x+1)+(x+2) solve for x

OpenStudy (anonymous):

can you do that or do you need help?

OpenStudy (anonymous):

HELP PLEAASEEEEE

OpenStudy (anonymous):

59 = x + (x + 1) So, 59 = 2x + 1, Solve for x, and the 2 numbers are x and x+1. Is this making sense?

OpenStudy (anonymous):

NOT REALLY I NEED AN EQUATION THAT MODELS THIS SITION AND THEN TO FIND THE TWO INTEGERS

OpenStudy (anonymous):

59=x+(x+1)+(x+2) this equation models the situation, then just solve for x to find the first integer, add 1 for the next integer, and add 2 to x to find the 3rd integer

OpenStudy (anonymous):

the simplified equation becomes 59=3x+3 can you solve it now?

OpenStudy (anonymous):

Well, the equation is what I put in my last post: 59 = x + (x + 1) Solving: 59 = 2x + 1 58 = 2x, so x = 29. So, your 2 integers are 29 and 30. Because 30 = x+1

OpenStudy (anonymous):

Also, the reason you have x and x+1 is that x+1 is the next integer after x.

OpenStudy (anonymous):

OK THANKS OTHER KID BUT TCARROLLO10 EXPLAINS IT BETTER THANKS

OpenStudy (anonymous):

I think @Studentc14 was thinking the problem asked for 3 integers.

OpenStudy (anonymous):

YEA ME TOO

OpenStudy (anonymous):

But your cool now!

OpenStudy (anonymous):

THANKS LOL

OpenStudy (anonymous):

oh yes, sorry i read that wrong...silly me

OpenStudy (anonymous):

np, and nice working with both of you.

OpenStudy (anonymous):

YOU TOO